The weather for today’s Clermont visit was a little humid and got a bit warm. We decided to start the May 28th visit at 9:00AM, an hour earlier than when we have been meeting.
All the colonies were examined for brood and most had brood in some or all phases – egg, larval, and capped. The level of swarming is not known so a few of the colonies that did not have evidence of queen activity may have swarmed and the new queen should be productive at our next visit. If needed, a nuc will be combined into any colony that is missing their queen. Most of the nucs had evidence of productive queens. Nucs not used for queen replacement will be moved into larger hives as the season progresses.
Several strong nucs can be used in re-queening as needed. One nuc that did not appear to have a viable queen was supplemented with two frames of brood and bees. These two frames also had several capped (or ripe) queen cells. The frame donor colony had additional queen cells that will likely be required if the colony swarms.
All residual Formic Pro™ mite treatment and mite drop boards were removed. Additional supers of foundation were added to several colonies in anticipation of continuing nectar flow. Top supers are starting to fill and have mostly capped honey. This past week, Locust trees and tulip popular have started to bloom. Abundant yellow Butter Cups are blooming in the fields surrounding the Clermont apiary.
Clermont continues to be a productive apiary with solid bee activity.